# DQ-DETR: DETR with Dynamic Query for Tiny Object Detection

### v2

@@ Abstract
Despite previous DETR-like methods having performed successfully in generic object detection, tiny object detection is still a challenging task for them since the positional information of object queries is not customized for detecting tiny objects, whose scale is extraordinarily smaller than general objects. Also, DETR-like methods using a fixed number of queries make them unsuitable for aerial datasets, which mostly contain tiny objects, and the numbers of instances are imbalanced between different images. Thus, we present a simple yet effective model, DQ-DETR, consisting of three components: categorical counting module, counting-guided feature enhancement, and dynamic query selection to solve the above-mentioned problems. DQ-DETR uses the prediction and density maps from the categorical counting module to dynamically adjust the number and positional information of object queries. Our model DQ-DETR outperforms previous CNN-based and DETR-like methods, achieving state-of-the-art mAP 30.2% on the AI-TOD-V2 dataset, which mostly consists of tiny objects.
Keywords DETR  $\cdot$ Tiny Object Detection
@@ 1 Introduction
Convolutional neural networks (CNNs) are superior to processing the RGB semantic and spatial texture features. Most object detection methods are primarily based on CNNs. For example, Faster R-CNN [1] introduces a region proposal network to generate potential object regions. FCOS [2] apply a center prediction branch to increase the quality of the bounding boxes.
@@
However, CNNs are unsuitable for attaining long-range dependency in the image, restricting the detection performance. Recently, DETR [3] incorporates CNN and transformer architecture to establish a new object detection framework. DETR utilizes the transformer encoder to integrate the partition image patches and pass them with the learnable object queries to the transformer decoder for final detection results. Moreover, a series of DETR-like methods [4, 5, 6, 7] aim to advance DETR performance and accelerate DETR convergence speed. For example, Deformable-DETR [4] uses multi-scale feature maps to improve its ability to detect different sizes of objects. Also, the use of deformable attention modules can not only capture more informative and contextually relevant features but accelerate training convergence as well.
@@
Table 1: Comparison of DETR-like models’ queries strategy under different situations.
Sparse Dense Imbalance Characteristics
Deformable DETR [4] Sparse Queries (K=300) with One-To-One Assignment; Low Recall
DDQ-DETR [8] Dense Distinct Queries (K=900); Low Recall if #Object $\gg$ #Query
DQ-DETR(Ours) Dynamically adjust the "Number" and "Position" of Queries
@@
We argue that the previous DETR-like methods are inappropriate in aerial images, which only contain tiny objects and have an imbalance of instances between different images. In the previous DETR-like methods, the object queries used in the transformer decoder do not consider the number and position of instances in the image. Generally, they apply the fixed number K of object queries, where K represents the maximum number of the detection objects, e.g., K=100, 900 in DETR and DINO-DETR, respectively. DETR [3] and Deformable-DETR [4] apply a fixed number of sparse queries, suffering a low recall rate. To address the problem, DDQ [8] select dense distinct queries, K=900, with a class-agnostic NMS based on a hand-designed IoU threshold. Though DDQ applies dense queries for detection, the number of queries is still limited.
@@
The fixed number of queries will cause low detection accuracy in aerial datasets since the number of objects may vary drastically in different images. In the AI-TOD-V2 dataset [9], some images have more than 1500 objects, but others have less than 10 objects. Under the situation that the number of objects in images is more than the DETR’s query number K, a low recall rate is an expected issue. Using smaller K restricts the recall of the objects in dense images, leaving lots of instances undetected (FN). While using a large K in the sparse images not only introduces many underlying false positive samples (FP) but also causes a waste of computing resources since the computing complexity in the decoder’s self-attention modules grows quadratic with the number of queries K.
@@
Besides, in the previous DETR-like methods, the object queries used in the transformer decoder do not consider the position of instances in the image. The position of object queries is a set of learned embeddings, which are irrelevant to the current image and do not have explicit physical meaning to tell where the queries are focusing on. The static positions of object queries are unsuitable for aerial datasets, where the distribution of instances varies extremely in different images, i.e., some images contain dense objects concentrated in specific areas, while some only have a few objects scattered throughout the images.
@@
Stemming on the above-mentioned weakness, we proposed a novel DETR-like method named DQ-DETR which mainly focuses on dynamically adapting the numbers of queries and enhancing the position of queries to locate the tiny objects precisely. In this work, we propose a dynamic query selection module for adaptively choosing different numbers of object queries in DETR’s decoder stage, resulting in fewer FP in sparse images and fewer FN in dense images. Moreover, we generate the density maps and estimate the number of instances in an image by the categorical counting module. The number of object queries is adjusted based on the predicted counting number. In addition, we aggregate the density maps with the visual feature from the transformer encoder to reinforce the foreground features, enhancing the spatial information for tiny objects. The strengthened visual feature will be further used to improve the positional information of object queries. As such, we can simultaneously handle the images with few and crowded tiny objects by dynamically adjusting the number and position of object queries used in the decoder.
In summary, our contributions are summarized as follows:
@@
    Experimental result shows that our proposed DQ-DETR significantly surpasses the state-of-the-art method by 16.6%, 20.5% in terms of $\mathrm{AP}$, $\mathrm{AP_{vt}}$ on the AI-TOD-V2 dataset.
@@ 2 Related work
2.1 DETR-like Methods.
DETR [3] proposes an end-to-end object detection framework based on the transformer named DETR (DEtection TRansformer), where the transformer encoder extracts instance-level features from an image, and the transformer decoder uses a set of learnable queries to probe and pool features from images. While DETR achieves comparable results with the previous classical CNN-based detectors [1, 2], it suffers severely from the problem of slow training convergence, needing 500 epochs of training to perform well. Many follow-up works have attempted to address the slow training convergence of DETR from different perspectives.
Some point out that the cause of DETR’s slow convergence is due to the unstable Hungarian matching and the cross-attention mechanism in the transformer decoder. [10] proposes an encoder-only DETR, discarding the transformer decoder. Dynamic DETR [11] designs an ROI-based dynamic attention mechanism in the decoder that can focus on regions of interest from a coarse-to-fine manner. Deformable-DETR [4] proposes an attention module that only attends to a few sampling points around a reference point. DN-DETR [5] introduces denoising training to reduce the difficulty of bipartite graph matching.
Another series of works makes improvements in decoder object queries. Since the object queries are just a set of learnable embedding in DETR [3],  [12, 13, 7] imputes the slow convergence of DETR to the implicit physical explanation of object queries. Conditional DETR [13] decouples the decoder’s cross-attention formulation and generates conditional queries based on reference coordinates. DAB-DETR [7] formulates the positional information of object queries as 4-D anchor boxes (x, y, w, h) that are used to provide ROI (Region of Interest) information for probing and pooling features.
2.2 Tiny Object Detection.
@@
Detecting small objects is challenging due to their lack of pixels. Early works apply data augmentation to oversample the instance of tiny objects. For example, [14], copy-paste small objects into the same image. [15] proposes a K sub-policies that automatically transform features from the instance level. In addition, several approaches, such as [9, 16, 17, 18] indicate that traditional Inter Section Union (IoU) metrics are ill-suited for tiny objects. When the object size difference is significant, IoU becomes highly sensitive. To design the appropriate metrics for the tiny object, DotD [18] considers the object’s absolute and relative size to formulate a new loss function. [16, 17, 18] design a new label assignment based on Gaussian distribution, which alleviates the sensitivity of the object size.
However, these methods highly rely on the predefined threshold, which is unstable to a different dataset. Instead, our DQ-DETR uses a counting-guided feature enhancement module to improve the visual features with more spatial information about tiny objects’ scale and location. Also, although the above-mentioned DETR-like methods have improved the formulation of queries, they are not specially designed to detect tiny objects. For example, the object queries in [7, 13, 3, 5] are learned from the training data and remain the same for different input images. Our proposed DQ-DETR is the first DETR-like model that focuses on tiny object detection. DQ-DETR dynamically adjusts the number of object queries and enhances the position information of queries in order to detect tiny objects precisely under imbalanced situations in aerial datasets.
@@ 3.1 Overview
The overall structure of DQ-DETR is shown in Fig. 1. As a DETR-like method, DQ-DETR is an end-to-end detector that contains a CNN backbone, a deformable transformer encoder [4], a deformable transformer decoder [4], and several prediction heads. We further implement a new categorical counting module, a counting-guided feature enhancement module, and dynamic query selection based on DETR’s architecture. Given an input image, we first extract multi-scale features with a CNN backbone and feed them into the transformer encoder to attain the encoder’s visual feature. Afterward, our categorical counting module determines how many object queries are used in the transformer decoder, as shown in Fig. 1(a). Besides, we propose a novel counting-guided feature enhancement module, shown in Fig. 1(b), to strengthen the encoder’s visual feature with spatial information of tiny objects. Last, the positional information of object queries will be refined via dynamic query selection, shown in Fig. 1(c). The following section will describe the proposed Categorical Counting module, Counting-Guided Feature Enhancement, and Dynamic Query Selection.
3.2 Reconstruction of Encoder’s Feature Map
@@
In our proposed categorical counting module, we will perform convolution operation on the transformer encoder features. Hence, we reconstruct the flattened encoder’s multi-scale visual features by reshaping its spatial dimension, resulting in 2-D feature maps ${S}_{i}\in\mathbb{R}^{d\times h_{i}\times w_{i}}$. We denote the reconstructed encoder’s multi-scale visual features as EMSV features for brevity.
@@ 3.3 Categorical Counting Module
We take the largest feature map $S_{1}$ of the EMSV features and generate the density map ${F}_{c}$ through the density extractor. The input feature map $S_{1}$ is firstly fed into a $1\times 1$ convolution layer for channel reduction ($\mathbb{R}^{b\times 256\times h\times w}\rightarrow\mathbb{R}^{b\times 512\times h\times w}$). Then, it is sent into a series of dilated convolution layers to acquire a density map ${F}_{c}$, which contains counting-related information. Specifically, dilated convolution layers enlarge the receptive field and capture rich long-range dependency for tiny objects.
@@
Lastly, we estimate the counting number $N$, i.e., the number of instances per image, by a classification head and categorize them into four levels, which are $N\leq 10$, $10<N\leq 100$, $100<N\leq 500$, and $N>500$. The classification head consists of two linear layers. Further, the numbers 10, 100, and 500 are selected based on the dataset’s characteristics, i.e., the mean and standard deviation of the number of instances $N$ per image. Notably, we do not use the regression head as in the traditional crowd-counting methods, which regresses the counting number to a specific number. We attribute the reason to the drastic difference of the number of instances in each image, where $N$ ranges from 1 to 2267 in different images of AI-TOD-V2. It is difficult to regress an accurate number, hurting the detection performance.
@@ 3.4 Counting-Guided Feature Enhancement Module (CGFE)
The EMSV feature will be refined with the density map from the categorical counting module through the proposed Counting-Guided Feature Enhancement Module (CGFE) to improve the spatial information of tiny objects. The refined features will further be used to enhance the position information of queries. This module comprises a spatial cross-attention and a channel attention operations.
@@
The generated spatial attention map ${W}_{s,i}$ multiplies with EMSV feature $S_{i}$ element-wisely and further obtain the spatial-intensified features ${E}_{i}$, as shown in Eq. 2.
@@ 3.5 Dynamic Query Selection
In dynamic query selection, we firstly use the classification result from the categorical counting module to determine the number of queries $K$ used in the transformer decoder. The four classification classes in the categorical counting module correspond to four distinct numbers of queries, which are $K$ = 300, 500, 900, and 1500, i.e., if the image is classified as $N\leq 10$, we will use $K=300$ queries in the further detection task, and so forth.
@@
For queries formulation, we follow the idea of DAB-DETR [7], where the queries are composed of content and positional information. The content of queries is a high-dimension vector, while the position of queries is formulated as a 4-D anchor box (x, y, w, h) to accelerate training convergence.
Further, we use the intensified multi-scale feature maps ${F}_{t}$ from the previous CGFE module to improve the content ${Q}_{content}$ and position ${Q}_{position}$ of queries. The each layer of ${F}_{t}$ are firstly flattened into pixel level and concatenate together, forming ${F}_{flat}\in\mathbb{R}^{b\times 256\times hw}$. The top-K features will be selected as priors to enhance decoder queries, where $K$ is the number of queries used in the transformer decoder stage. The selection is based on the classification score. We feed ${F}_{flat}$ into an FFN for the object classification task and generates the classification score $\in\mathbb{R}^{b\times m\times hw}$, where m is the number of object classes in the dataset. Afterwards, we will generate the content and position of queries using the selected top-K features ${F}_{select}$.
@@
Since the features ${F}_{select}$ are selected from ${F}_{t}$, which is generated from the previous CGFE module, they contain abundant scale and location information of tiny objects. Hence, the enhanced content and position of object queries are tailored based on each image’s crowded or sparse situation and make the queries easier to localize the tiny objects in transformer decoder stage.

### v3

@@ Abstract
Despite previous DETR-like methods having performed successfully in generic object detection, tiny object detection is still a challenging task for them since the positional information of object queries is not customized for detecting tiny objects, whose scale is extraordinarily smaller than general objects. Additionally, the fixed number of queries used in DETR-like methods makes them unsuitable for detection if the number of instances is imbalanced between different images. Thus, we present a simple yet effective model, DQ-DETR, consisting of three components: categorical counting module, counting-guided feature enhancement, and dynamic query selection to solve the above-mentioned problems. DQ-DETR uses the prediction and density maps from the categorical counting module to dynamically adjust the number and positional information of object queries. Our model DQ-DETR outperforms previous CNN-based and DETR-like methods, achieving state-of-the-art mAP 30.2% on the AI-TOD-V2 dataset, which mostly consists of tiny objects. Our code will be available at https://github.com/Katie0723/DQ-DETR.
Keywords:
Detection Transformer Query Selection Tiny Object Detection
@@ 1 Introduction
Convolutional neural networks (CNNs) excel at processing the RGB semantic and spatial texture features. Most object detection methods are primarily based on CNNs. For example, Faster R-CNN [16] introduces a region proposal network to generate potential object regions. FCOS [19] applies a center prediction branch to increase the quality of the bounding boxes.
@@
However, CNNs are unsuitable for capturing long-range dependencies in the image, restricting the detection performance. Recently, DETR [2] incorporates CNN and transformer architecture to establish a new object detection framework. DETR utilizes the transformer encoder to integrate the partitioned image patches and passes them with the learnable object queries to the transformer decoder for final detection results. Moreover, a series of DETR-like methods [30, 7, 27, 11] aim to advance DETR performance and accelerate convergence speed. For example, Deformable-DETR [30] uses multi-scale feature maps to improve its ability to detect different sizes of objects. Also, the use of deformable attention modules can not only capture more informative and contextually relevant features but accelerate training convergence as well.
@@
Table 1: Comparison of DETR-like models’ query strategies under different situations.
{tblr}
hlines, vline2-5 = -, column2-4 = c, column5 = c, & Sparse Dense Imbalance Characteristics
Deformable DETR [30] ✓ Sparse Queries (K=300) with One-To-One Assignment; Low Recall
DDQ-DETR [28] ✓ ✓ Dense Distinct Queries (K=900); Low Recall if #Object $\gg$ #Query
DQ-DETR (Ours) ✓ ✓ ✓ Dynamically adjust the "Number" and "Position" of Queries
@@
In this work, we argue that the previous DETR-like methods are inappropriate in aerial image datasets, which only contain tiny objects and have an imbalance of instances between different images. In the previous DETR-like methods, the object queries used in the transformer decoder do not consider the number and position of instances in the image. Generally, they apply a fixed number K of object queries, where K represents the maximum number of the detection objects, e.g., K=100, 300 in DETR and Deformable-DETR, respectively. DETR [2] and Deformable-DETR [30] apply a fixed number of sparse queries, suffering from a low recall rate. To address this problem, DDQ [28] selects dense distinct queries, K=900, with a class-agnostic NMS based on a hand-designed IoU threshold. Though DDQ applies dense queries for detection, the number of queries is still limited.
@@
However, aerial datasets often exhibit imbalances in the distribution of instances across different images. A fixed number of queries can lead to poor detection accuracy when the number of objects varies drastically between images. For example, in the AI-TOD-V2 dataset [24], some images have more than 1500 objects, but others have less than 10 objects. Under the situation that the number of objects in images is more than DETR’s query number K, a low recall rate is an expected issue. Using smaller K restricts the recall of the objects in dense images, leaving many instances undetected (FN). Conversely, using a large K in the sparse images not only introduces many underlying false positive samples (FP) but also causes a waste of computing resources since the computing complexity in the decoder’s self-attention layers grows quadratic with the number of queries K.
@@
Furthermore, in the previous DETR-like methods, the object queries do not consider the position of instances in the image. The position of object queries is a set of learned embeddings, which are irrelevant to the current image and do not have explicit physical meaning to tell where the queries are focusing on. The static positions of object queries are unsuitable for aerial image datasets, where the distribution of instances varies extremely in different images, i.e., some images contain dense objects concentrated in specific areas, while some only have a few objects scattered throughout the images.
@@
Stemming from the above-mentioned weakness, we propose a novel DETR-like method named DQ-DETR, which mainly focuses on dynamically adapting the numbers of queries and enhancing the position of queries to locate the tiny objects precisely. In this study, we propose a dynamic query selection module for adaptively choosing different numbers of object queries in DETR’s decoder stage, resulting in fewer FP in sparse images and fewer FN in dense images. Moreover, we generate the density maps and estimate the number of instances in an image by the categorical counting module. The number of object queries is adjusted based on the predicted counting number. In addition, we aggregate the density maps with the visual feature from the transformer encoder to reinforce the foreground features, enhancing the spatial information for tiny objects. The strengthened visual feature will be further used to improve the positional information of object queries. As such, we can simultaneously handle the images with few and crowded tiny objects by dynamically adjusting the number and position of object queries used in the decoder.
Our contributions are summarized as follows:
@@
    Experimental result shows that our proposed DQ-DETR significantly surpasses the state-of-the-art method by 16.6%, 20.5% in terms of AP, $\text{AP}_{vt}$ on the AI-TOD-V2 dataset.
@@ 2 Related work
2.0.1 Tiny Object Detection.
@@
Detecting small objects is challenging due to their lack of pixels. Early works apply data augmentation to oversample the instance of tiny objects. For example, [6], copy-paste small objects into the same image. [31] proposes a K sub-policies that automatically transform features from the instance level. In addition, several approaches, such as [24, 20, 25, 26], indicate that traditional Intersection over Union (IoU) metrics are ill-suited for tiny objects. When the object size difference is significant, IoU becomes highly sensitive. To design the appropriate metrics for the tiny object, DotD [26] considers the object’s absolute and relative size to formulate a new loss function. [20, 25, 26] design a new label assignment based on Gaussian distribution, which alleviates the sensitivity of the object size. However, these methods heavily rely on the predefined threshold, which is unstable for a different dataset.
2.0.2 DETR-like Methods.
DETR [2] proposes an end-to-end object detection framework based on the transformer, where the transformer encoder extracts instance-level features from an image, and the transformer decoder uses a set of learnable queries to probe and pool features from images. While DETR achieves comparable results with the previous classical CNN-based detectors [16, 19], it suffers severely from the problem of slow training convergence, needing 500 epochs of training to perform well. Many follow-up works have attempted to address the slow training convergence of DETR from different perspectives.
Some argue that DETR’s slow convergence stems from the instability of Hungarian matching and the cross-attention mechanism in the transformer decoder. [18] proposes an encoder-only DETR, discarding the transformer decoder. Dynamic DETR [3] designs an ROI-based dynamic attention mechanism in the decoder that can focus on regions of interest from a coarse-to-fine manner. Deformable-DETR [30] proposes an attention module that only attends to a few sampling points around a reference point. DN-DETR [7] introduces denoising training to reduce the difficulty of bipartite graph matching.
Another series of works makes improvements in decoder object queries. Since the object queries are just a set of learnable embedding in DETR,  [22, 13, 11] imputes the slow convergence of DETR to the implicit physical explanation of object queries. Conditional DETR [13] decouples the decoder’s cross-attention formulation and generates conditional queries based on reference coordinates. DAB-DETR [11] formulates the positional information of object queries as 4-D anchor boxes (x, y, w, h) that are used to provide ROI (Region of Interest) information for probing and pooling features. Although DETR-like methods have improved the formulation of queries, they are constrained in their ability to handle tiny objects and datasets with widely varying numbers of objects. The object queries in these methods are learned from the training data and the number of queries remains the same across different input images.
Our proposed DQ-DETR stands out as the first DETR-like model specifically designed to detect tiny objects and dynamically adjust the number of queries to enhance precision in imbalanced datasets.
@@ 3.1 Overview
The overall structure of DQ-DETR is shown in Fig. 1. As a DETR-like method, DQ-DETR is an end-to-end detector that contains a CNN backbone, a deformable encoder and decoder [30], and several prediction heads. We further implement a new categorical counting module, a counting-guided feature enhancement module, and dynamic query selection based on DETR’s architecture. Given an input image, we first extract multi-scale features with a CNN backbone and feed them into the transformer encoder to attain visual features. Afterward, our categorical counting module determines how many object queries are used in the transformer decoder, as shown in Fig. 1(a). Besides, we propose a novel counting-guided feature enhancement module, as illustrated in Fig. 1(b), to strengthen the encoder’s visual features with spatial information of tiny objects. Last, the object queries are refined with additional information about the location and size of tiny objects through dynamic query selection, as shown in Fig. 1(c). The following section will describe the proposed Categorical Counting Module, Counting-Guided Feature Enhancement, and Dynamic Query Selection.
3.2 Unflattening of Encoder’s Feature Map
@@
In our proposed categorical counting module, we apply dilated convolution operations on the transformer encoder features. Hence, we unflatten the encoder’s multi-scale visual features by reshaping its spatial dimension, resulting in 2-D feature maps ${S}_{i}\in\mathbb{R}^{d\times h_{i}\times w_{i}}$. We denote the reconstructed encoder’s multi-scale visual features as EMSV features for brevity.
@@ 3.3 Categorical Counting Module
We take the largest feature map $S_{1}$ of the EMSV features and generate the density map ${F}_{c}$ through the density extractor. The input feature map $S_{1}$ is sent into a series of dilated convolution layers to acquire a density map ${F}_{c}$, which contains counting-related information. Specifically, dilated convolution layers enlarge the receptive field and capture rich long-range dependency for tiny objects.
@@
Lastly, we estimate the counting number $N$, i.e., the number of instances per image, by a classification head and categorize them into four levels, which are $N\leq 10$, $10<N\leq 100$, $100<N\leq 500$, and $N>500$. The classification head consists of two linear layers. Further, the numbers 10, 100, and 500 are selected based on the AI-TOD-V2 dataset’s characteristics, i.e., the mean and standard deviation of the number of instances $N$ per image. Notably, we do not use the regression head as in the traditional crowd-counting methods, which regresses the counting number to a specific number. We attribute the reason to the drastic difference in the number of instances in each image, where $N$ ranges from 1 to 2267 in different images of AI-TOD-V2. It is difficult to regress an accurate number, hurting the detection performance.
@@ 3.4 Counting-Guided Feature Enhancement Module (CGFE)
The EMSV feature is refined using the density map from the categorical counting module through the proposed Counting-Guided Feature Enhancement Module (CGFE) to improve the spatial information of tiny objects. The refined features are then used to enhance the position information of queries.
This module comprises spatial cross-attention and channel attention operations [23].
@@
The generated spatial attention map ${W}_{s,i}$ multiplies with EMSV feature $S_{i}$ element-wisely and further obtains the spatial-intensified features ${E}_{i}$, as shown in Eq. 2.
@@ 3.5 Dynamic Query Selection
In dynamic query selection, we first use the classification result from the categorical counting module to determine the number of queries $K$ used in the transformer decoder. The four classification classes in the categorical counting module correspond to four distinct numbers of queries, which are $K$ = 300, 500, 900, and 1500, i.e., if the image is classified as $N\leq 10$, we use $K=300$ queries in the subsequent detection task, and so forth.
@@
For query formulation, we follow the idea of DAB-DETR [11], where the queries are composed of content and positional information. The content of queries is a high-dimension vector, while the position of queries is formulated as a 4-D anchor box (x, y, w, h) to accelerate training convergence.
Further, we use the intensified multi-scale feature maps ${F}_{t}$ from the previous CGFE module to improve the content ${Q}_{content}$ and position ${Q}_{position}$ of queries. Each layer of ${F}_{t}$ is firstly flattened into pixel level and concatenated together, forming ${F}_{flat}\in\mathbb{R}^{b\times 256\times hw}$. The top-K features will be selected as priors to enhance decoder queries, where $K$ is the number of queries used in the transformer decoder stage. The selection is based on the classification score. We feed ${F}_{flat}$ into an FFN for the object classification task and generate the classification score $\in\mathbb{R}^{b\times m\times hw}$, where m is the number of object classes in the dataset. Consequently, we generate the content and position of queries using the selected top-K features ${F}_{select}$.
@@
Since the features ${F}_{select}$ are selected from ${F}_{t}$, which is generated from the previous CGFE module, they contain abundant scale and location information of tiny objects. Hence, the enhanced content and position of object queries are tailored based on each image’s density (crowded or sparse), facilitating easier localization of tiny objects in the transformer decoder.